# Locally Private Histograms in All Privacy Regimes

### v1

@@ Abstract
Frequency estimation, a.k.a. histograms, is a workhorse of data analysis, and as such has been thoroughly studied under differentially privacy. In particular, computing histograms in the local model of privacy has been the focus of a fruitful recent line of work, and various algorithms have been proposed, achieving the order-optimal $\ell_{\infty}$ error in the high-privacy (small $\varepsilon$) regime while balancing other considerations such as time- and communication-efficiency. However, to the best of our knowledge, the picture is much less clear when it comes to the medium- or low-privacy regime (large $\varepsilon$), despite its increased relevance in practice. In this paper, we investigate locally private histograms, and the very related distribution learning task, in this medium-to-low privacy regime, and establish near-tight (and somewhat unexpected) bounds on the $\ell_{\infty}$ error achievable. Our theoretical findings emerge from a novel analysis, which appears to improve bounds across the board for the locally private histogram problem. We back our theoretical findings by an empirical comparison of existing algorithms in all privacy regimes, to assess their typical performance and behaviour beyond the worst-case setting.
@@ 1.1 Prior work
RAPPOR $\checkmark$ $k$ $\sqrt{\frac{\log k}{n\min(1,\varepsilon^{2})}}$ $\frac{kne^{\varepsilon/2}}{\left(e^{\varepsilon/2}-1\right)^{2}}$
@@
RHR ([CKÖ23, Theorem 3.1]) $\times$ $\ell$ $\frac{k}{n\min\left(\left(e^{\varepsilon/2}-1\right)^{2},e^{\varepsilon},2^{\ell},k\right)}$
No name ([CKÖ23, Theorem 3.1]) $\times$ $\ell$ $\sqrt{\frac{\log k}{n\min(\varepsilon,\varepsilon^{2},\ell)}}$ $\frac{ke^{2\varepsilon/\ell}}{n\ell\left(e^{\varepsilon/\ell}-1\right)^{2}}$ ()($\dagger$)( † )
@@ 1.2 Overview of results
After providing some background and setting notation, we establish our theoretical results (upper bounds on the error rate) in Section 3, followed by information-theoretic lower bounds in Section 4. Section 5 then contains our empirical findings; finally, we discuss our results and potential future work in Section 6. Some miscellaneous proofs, omitted for clarity of exposition, can be found in the appendix.
@@ 3.4 Upper bound for Projective Geometry Response
Projective geometry response (PGR), introduced by [FNNT22] achieves optimal rates for $\ell_{2}^{2}$ error, communication, and near-optimal processing time. The protocol is based on the general template established in [ASZ19]: specifically, PGR relies on a set structure defined by projective planes, as detailed next. For $d$ a prime power and $k=\frac{d^{t}-1}{d-1}$, the authors define a $t$-dimensional vector space $\mathbb{F}^{t}_{d}$, where each element $x\in[k]$ is represented by one of the canonical basis vectors. These basis vectors in turn each uniquely determine a projective plane $S(x)$, such that there are $|S(x)|=\frac{d^{t-1}-1}{d-1}$ “high probability” elements. Every one of these sets in turn has an intersection with every other set of size $|S(x)\cap S(x^{\prime})|=\frac{d^{t-2}-1}{d-1}$. By choosing a prime power $d\approx e^{\varepsilon}+1$, optimal error is achieved.
@@
Before establishing the theorem, we briefly recall the details of PGR as described in [FNNT22], as well as the setting of its parameters. The protocol follows the structure established in [ASZ19] which associates each input $x\in[k]$ with a set of high probability outputs $S(x)$, all of the same cardinality. Denote the size of each subset as
$s=|S(x)|=\frac{d^{t-1}-1}{d-1}$
and the size of the intersection of any two subsets corresponding to distinct $x,x^{\prime}$ as
$c=|S(x)\cap S(x^{\prime})|=\frac{d^{t-2}-1}{d-1}\,,$
and note that $s-c=d^{t-2}$. The probability of returning any element $y$ from the output alphabet, given $x$ as an input is,
@@
(While we could instead choose $d<e^{\varepsilon}+1$ and set the inequality to be an undershoot by a factor two, it seems that this direction leads to a better error bound.)
@@
Note that, as introduced above, we must have $k=\frac{d^{t}-1}{d-1}$ for a prime power $d$ satisfying (11). Other values of $k$ must be rounded up, losing up to a factor $O(e^{\varepsilon})$ in the domain size (i.e., working instead with a domain size $k^{\prime}=O(e^{\varepsilon}k)$). We hereafter ignore this detail, which does not affect the final bound of Theorem 8 unless $\varepsilon\gg\log k$; and assume that $k$ is of the form $k=\frac{d^{t}-1}{d-1}$. For the same reason, we additionally can assume $t\geq 3$.
@@
and recall that $s-c=d^{t-2}$. So using Eq. 12
@@
where the final inequality comes from the fact that when $d\geq 2$ this is at most the sum of the geometric series $\sum_{i=0}^{\infty}1/2^{i}=2$. For $k$ we have a similar series, with the addition of a single $d$ term:
@@
$\mathbb{E}\left[\lVert\hat{q}-q{\rVert}_{\infty}\right]\in{O\left(\sqrt{\frac{\log k}{ne^{\varepsilon}}}+\frac{\log k}{n\varepsilon}\log n\right)}\,,$
@@ 5 Empirical evaluation and findings
Figure 2: Log–log plot of median $\ell_{\infty}$ error with upper and lower quartiles, by $\varepsilon$. Lower and upper bounds discussed in this work included for comparison. No normalisation or clipping has been applied leading to $\ell_{\infty}>2$ in the high privacy regime. Behaviour when $\varepsilon$ is large is likely caused by floating-point precision errors.
Figure 2 demonstrates RAPPOR and Subset Selections’ performance against its theoretical upper and lower bounds: we suspect the tail behaviour is due to finite precision randomness when probabilities become close to 1 (or 0 respectively), [^9] Unfortunately the libraries used did not include a additionally one can see subset selection become worse in the regime where $\varepsilon>\log k$, until the subset size reaches 1 and it becomes practically non-private frequency estimation. We can see a minor transition in the low-privacy regime to the Local Glivenko–Cantelli behaviour, which is much more apparent for large $k$ as while simple analysis of the sub–Gaussian parameters begins to diverge from the lower bound. Figure 3 demonstrates this more clearly. All of the experiments together took approximately 4 hours on a 2022 M2 Macbook air.
@@ 6 Discussion and future work
Tighter analysis of other asymmetric protocols.
While we analyse RAPPOR, which admits a simple analysis due to the independence of coordinates, and Projective Geometry Response which represents the state of the art in low-communication LDP protocols, we believe the tools introduced in this paper are applicable to all LDP protocols. Of particular interest would be Subset Selection, which has optimal mutual information between inputs and outputs of the local randomiser, and the two Hadamard Response protocols that we include in our empirical analysis.
@@
As mentioned in the introduction, one of the key motivations for studying locally private histogram estimation in the low privacy regime is the implication for histogram estimation in the shuffle model of privacy (see, e.g.[GKMP20, GGK+21, BC20, CZ22]), in light of the “plug-and-play” amplification-by-shuffling results allowing to “translate” the former into the latter (cf., for instance, [FMT21]). However, the results obtained in this work, both upper and lower bound, seem to indicate this approach would either not lead to the optimal result, or be far from “plug-and-play” and instead require some non-trivial analysis. Indeed, at a high level, obtaining optimal error rates for $\ell_{\infty}$ frequency estimation (or distribution learning) via a naive blackbox application of amplification-by-shuffling seems to require the LDP protocol we start with to have an error dependence decaying, for large $\varepsilon$, as $\sqrt{{\log k}/(ne^{\varepsilon})}$: our results (both upper and lower bounds) indicate that this is not an achievable rate.

### v2

@@ Abstract
Frequency estimation, a.k.a. histograms, is a workhorse of data analysis, and as such has been thoroughly studied under differentially privacy. In particular, computing histograms in the local model of privacy has been the focus of a fruitful recent line of work, and various algorithms have been proposed, achieving the order-optimal $\ell_{\infty}$ error in the high-privacy (small $\varepsilon$) regime while balancing other considerations such as time- and communication-efficiency. However, to the best of our knowledge, the picture is much less clear when it comes to the medium- or low-privacy regime (large $\varepsilon$), despite its increased relevance in practice. In this paper, we investigate locally private histograms, and the very related distribution learning task, in this medium-to-low privacy regime, and establish near-tight (and somewhat unexpected) bounds on the $\ell_{\infty}$ error achievable. As a direct corollary of our results, we obtain a protocol for histograms in the shuffle model of differential privacy, with accuracy matching previous algorithms but significantly better message and communication complexity.
Our theoretical findings emerge from a novel analysis, which appears to improve bounds across the board for the locally private histogram problem. We back our theoretical findings by an empirical comparison of existing algorithms in all privacy regimes, to assess their typical performance and behaviour beyond the worst-case setting.
@@ 1.1 Prior work
RAPPOR $\checkmark$ $k$ $\sqrt{\frac{\log k}{n\min(1,\varepsilon^{2})}}$ $\frac{ke^{\varepsilon/2}}{n\left(e^{\varepsilon/2}-1\right)^{2}}$
@@
RHR [CKÖ23, Theorem 3.1] $\times$ $\ell$ $\frac{k}{n\min\left(\left(e^{\varepsilon/2}-1\right)^{2},e^{\varepsilon},2^{\ell},k\right)}$
No name [CKÖ23, Theorem 3.1] $\times$ $\ell$ $\sqrt{\frac{\log k}{n\min(\varepsilon,\varepsilon^{2},\ell)}}$ $\frac{ke^{2\varepsilon/\ell}}{n\ell\left(e^{\varepsilon/\ell}-1\right)^{2}}$ ()($\dagger$)( † )
@@
Histograms have also seen detailed analysis in the shuffle model as both a practical question [GGK+21, GKMP20, CSU+19, CZ22, BEM+17] and as a benchmark for reasoning about the power of the shuffle model [BC20a]. As two points of comparison we highlight [GGK+21] which introduced a multiple–message protocol for shuffled histograms achieving expected $\ell_{\infty}$-error $O({\sqrt{\log k\log(1/\delta)}/(\varepsilon n)})$ with $O({k^{1/100}})$ rounds of communication, and [BC20a] which demonstrated expected maximum error of $O(\log(1/\delta)/(\varepsilon^{2}n))$ with $k+1$ rounds. While our results do not achieve the latter bound, we show surprisingly that the former can be achieved in some fairly permissive parameter regimes with only a single message.
@@ 1.2 Overview of results
Implications for shuffle privacy.
By combining our new results on LDP protocols for histograms in the low-privacy regime with known “amplification-by-shuffling” results, we are able to obtain a simple, robust shuffle (approximate) DP protocol using only one message per user of logarithmic length, while achieving state-of-the-art:
Theorem 5 (Informal; see Theorem 12).
Shuffling the Projective Geometry Response protocol achieves expected $\ell_{\infty}$ error
${O\left(\frac{\sqrt{\log(k)\log(1/\delta)}}{n\varepsilon}\right)}$
for all $\varepsilon\in[\Omega(1/\sqrt{n}),1]$, with $O(\log k)$ bits of communication and a single message per user.
This matches that of the best known shuffle DP protocols, but with a much lower message and communication complexity. It was, to the best of our knowledge, still open whether achieving these “best of three worlds” guarantees was possible.
LDP protocols beyond worst-case bounds.
@@
After providing some background and setting notation, we establish our theoretical results (upper bounds on the error rate) in Section 3, followed by information-theoretic lower bounds in Section 4 and applications to shuffle privacy in Section 5. Section 6 then contains our empirical findings; finally, we discuss our results and potential future work in Section 7. Some miscellaneous proofs, omitted for clarity of exposition, can be found in the appendix.
@@ 3.4 Upper bound for Projective Geometry Response
Projective geometry response (PGR), introduced by [FNNT22] achieves optimal rates for $\ell_{2}^{2}$ error, communication, and near-optimal processing time. The protocol is based on the general template established in [ASZ19]: specifically, PGR relies on a set structure defined by projective planes, as detailed next. For a prime power $d$ and $k=\frac{d^{t}-1}{d-1}$, the authors define a $t$-dimensional vector space $\mathbb{F}^{t}_{d}$, where each element $x\in[k]$ is represented by one of the canonical basis vectors. These basis vectors in turn each uniquely determine a projective plane $S(x)$, such that there are $s=|S(x)|=\frac{d^{t-1}-1}{d-1}$ “high probability” elements. Every one of these sets in turn has an intersection with every other set of size $c=|S(x)\cap S(x^{\prime})|=\frac{d^{t-2}-1}{d-1}$. By choosing a prime power $d\approx e^{\varepsilon}+1$, optimal error is achieved.
@@
We briefly recall some details of PGR as described in [FNNT22]. We will make use of the size of each subset $s$ and the size of each intersection $c$ as described in the introduction to this section, as well as the probability of returning any element $y$ from the output alphabet, given $x$ as an input:
@@
(While we could instead choose $d<e^{\varepsilon}+1$ and set the inequality to be an undershoot by a factor two, an investigation in that direction led to worse constants and trickier analysis.)
@@
Note that, as introduced above, we must have $k=\frac{d^{t}-1}{d-1}$ for a prime power $d$ satisfying (11). Other values of $k$ must be rounded up, losing up to a factor $O(e^{\varepsilon})$ in the domain size (i.e., working instead with a domain size $k^{\prime}=O(e^{\varepsilon}k)$). We hereafter ignore this detail, which does not affect the final bound of Theorem 9 unless $\varepsilon\gg\log k$; and assume that $k$ is of the form stated above. For the same reason, we additionally can assume $t\geq 3$.
@@
and note that by simple application of Eq. 12 we get $s-c=d^{t-2}$. Applying the identity again we get:
@@
where the final inequality comes from the fact that with $d\geq 2$ this is bounded by the geometric series. For $k$ we have the same series, with the addition of a single $d$ term.
@@
$\mathbb{E}\left[\lVert\hat{q}-q{\rVert}_{\infty}\right]\in{O\left(\sqrt{\frac{\log k}{ne^{\varepsilon}}}+\frac{\log k}{n\varepsilon}\log n\right)}\,$
@@ 4 Worst-case, information-theoretic lower bounds
5 Amplification by shuffling
As mentioned in the introduction, one of the key motivations for studying locally private histogram estimation in the low privacy regime is the implication for histogram estimation in the shuffle model of privacy (see, e.g.[GKMP20, GGK+21, BC20b, CZ22]), in light of the “plug-and-play” amplification-by-shuffling results allowing to “translate” the former into the latter. Specifically, we will use the following result of Feldman, McMillan, and Talwar:
Theorem 11 ([FMT21, Theorem 3.1]).
For any domain $\mathcal{X}$, let $R\colon\mathcal{X}\to\mathcal{Y}$ be an $\varepsilon_{L}$-DP local randomiser; and let $S$ be the algorithm that given a tuple of $n$ messages $\vec{y}\in\mathcal{Y}^{n}$, samples a uniform random permutation $\pi$ over $[n]$ and outputs $(\vec{y}_{\pi(1)},\dots,\vec{y}_{\pi(n)})$. Then for any $\delta\in(0,1]$ such that $\varepsilon_{L}\leq\log\frac{n}{16\log(2/\delta)}$, $S\circ R^{n}$ is $(\varepsilon,\delta)$-DP, where
$\varepsilon\leq\log\left(1+8\frac{e^{\varepsilon_{L}}-1}{e^{\varepsilon_{L}}+1}\left(\sqrt{\frac{e^{\varepsilon_{L}}\log(4/\delta)}{n}}+\frac{e^{\varepsilon_{L}}}{n}\right)\right)\,.$
In particular, if $\varepsilon_{L}\geq 1$ then $\varepsilon=O\left(\sqrt{e^{\varepsilon_{L}}\log(1/\delta)/n}\right)$, and if $\varepsilon_{L}<1$ then $\varepsilon=O\left(\varepsilon_{L}\sqrt{\log(1/\delta)/n}\right)$.
This implies the following:
Lemma 7 (Amplification by shuffling).
Fix any $\delta\in(0,1]$, $\varepsilon\in(0,1]$, and $n$ such that $\varepsilon>16\sqrt{\log(4/\delta)/n}$. Then, for
$\varepsilon_{L}\coloneqq\log\frac{\varepsilon^{2}n}{256\log(4/\delta)}={\Theta\left(\log\frac{\varepsilon^{2}n}{\log(1/\delta)}\right)},$
shuffling the messages of $n$ users using the same $\varepsilon_{L}$-LDP randomizer satisfies (robust) $(\varepsilon,\delta)$-shuffle differential privacy.
Proof.
Note that for $\varepsilon,\delta$ as in the statement and $\varepsilon_{L}$ as defined, we have $0<\varepsilon_{L}\leq\log\frac{n}{16\log(2/\delta)}$. Applying Theorem 11, we get $(\varepsilon^{\prime},\delta)$ privacy for
$\varepsilon^{\prime}=\log\Big{(}1+8\underbrace{\frac{e^{\varepsilon_{L}}-1}{e^{\varepsilon_{L}}+1}}_{\leq 1}\Big{(}\frac{\varepsilon}{16}+\underbrace{\frac{\varepsilon^{2}}{256\log(4/\delta)}}_{\leq\varepsilon/16}\Big{)}\Big{)}\leq\varepsilon$
which proves the statement. ∎
Theorem 12.
For $n={\Omega\left(\frac{\log(1/\delta)}{\varepsilon^{2}}\right)}$ and $\varepsilon\in(0,1]$, Shuffled Projective Geometry Response achieves maximum error ${O\left(\frac{\sqrt{\log(k)\log(1/\delta)}}{n\varepsilon}\right)}$, with $O(\log k)$ bits of communication (and one single message) per user.
Proof.
For $n\geq\frac{500}{\varepsilon^{2}}\log\frac{4}{\delta}$, the restriction on $\varepsilon$ from Lemma 7 holds, and, setting $\varepsilon_{L}$ as in the lemma, we also have $\varepsilon_{L}=\Omega(1)$. We then invoke the bound of Theorem 9, focusing on first part of the bound, which dominates when $\frac{\log k}{n}\cdot\log^{2}n\leq\frac{\varepsilon_{L}^{2}}{e^{\varepsilon_{L}}}$. This leads to an upper bound on the error of the order
$\sqrt{\frac{\log k}{ne^{\varepsilon_{L}}}}$ $=\sqrt{\frac{\log(k)}{n\varepsilon^{2}\frac{n}{\log(1/\delta)}}}=\frac{\sqrt{\log(k)\log(1/\delta)}}{n\varepsilon}$
as desired. It only remains to argue that the first term of the bound did, indeed, dominate the error. As mentioned above, this is the case whenever
$\frac{e^{\varepsilon_{L}}}{\varepsilon_{L}^{2}}\cdot\log k\ll\frac{n}{\log^{2}n}$
for which a weaker, sufficient condition is $\frac{e^{\varepsilon_{L}}}{\varepsilon_{L}^{2}}\ll\frac{n}{\log^{2}n}$, that is, $n\gg e^{\varepsilon_{L}}$. But this follows from our setting of $\varepsilon_{L}$, such that $e^{\varepsilon_{L}}=\underbrace{\tfrac{\varepsilon^{2}}{256\log(4/\delta)}}_{\ll 1}\cdot n$. ∎
Comparing this result with the summary of local, shuffle, and central histogram error bounds available in [CZ22, Table 1] shows that with shuffled PGR achieves the best error of any protocol which sends a constant number of messages.
More specifically, focusing on 3 representative known protocols: the only known one-message-per-user protocol, due to [CZ22], achieves much worse error, [^9] And has the same restriction $n=\Omega(\log(1/\delta)/\varepsilon^{2})$ on the parameters. and requires $k$ bits of communication per user; while the protocol of [GGK+21], which achieves the same error as Theorem 12, uses $k^{\Omega(1)}$ $(\log k)$-bit messages per user. Finally, a protocol of [BC20a] does achieve better error, but at the cost of performing $k+1$ rounds, each with $\log k$ communication per user. Thus, our bounds demonstrates that shuffled PGR achieve state-of-the-art $\ell_{\infty}$–error with only one message of $\log k$ bits.
@@ 6 Empirical evaluation and findings
All of the experiments together took approximately 4 hours on a 2022 M2 Macbook air.
@@ 7 Discussion and future work
The logarithmic factor in the upper bound.
We prove our upper bounds using the result of [CK23], which is where the $\log n$ term appears. The authors conjectured that this term could be removed in general, but follow–up work [BV24] demonstrated by counter–example that this is not the case. They show that there is a distribution–dependent interpolation between the term being $T(n)/n$ and $T(n)\log n/n$. This logarithmic factor has implications for understanding which error regime is dominating, given the parameters of the algorithm $(\varepsilon,k,n)$. While it is possible to imagine that a careful application of the tools in [BV24] could resolve this matter in general or for a specific use–case, in the meantime we emphasise that empirical analysis is still crucial.
Tighter analysis of other protocols.
While we analyse RAPPOR which admits a simple analysis due to the independence of coordinates, and Projective Geometry Response which represents the state of the art in low-communication LDP protocols, we believe the tools introduced in this paper are applicable in a very general way to most LDP protocols. Of particular interest would be Subset Selection, which has optimal mutual information between inputs and outputs of the local randomiser, and the two Hadamard Response protocols that we include in our empirical analysis.
@@
While our shuffle DP result (Theorem 12) yields better error, communication, and number of messages, it does have a limitation on the parameter range, namely $n=\Omega(\log(1/\delta)/\varepsilon^{2})$. It would be interesting to weaken this requirement to match the central DP one (where the dependence on $\varepsilon$ is only linear). Moreover, one could hope to further improve the resulting error to the central DP bound (i.e., a $\min(\log k,\log(1/\delta))$ dependence instead of $\sqrt{\log k\log(1/\delta)}$), or, alternatively, prove a matching lower bound separating the two models.
Broader applicability of lower bound techniques.
While we prove lower bounds for the specific case of $\ell_{\infty}$ frequency estimation, the tools used are extremely general and we imagine that their application could provide new lower bounds for a variety of problems in the LDP setting, especially with the low–privacy regime in mind.
Acknowledgments.
The authors would like to thank Guy Blanc for the proof of Lemma 2, and Albert Cheu for insightful discussions regarding the use of amplification by shuffling (Lemma 7).
@@ Appendix A Deferred miscellaneous proofs
Appendix B More empirical analysis and bounds plots
Figure 3: Log–log plot of median $\ell_{\infty}$ error with upper and lower quartiles, by $\varepsilon$. Lower and upper bounds discussed in this work included for comparison. No normalisation or clipping has been applied leading to $\ell_{\infty}>2$ in the high privacy regime.
Figure 3 demonstrates RAPPOR and Subset Selections’ performance against its theoretical upper and lower bounds. One can see subset selection become worse in the regime where $\varepsilon>\log k$, until the subset size reaches 1 and it becomes practically non-private frequency estimation. We can see a minor transition in the low-privacy regime to the Local Glivenko–Cantelli behaviour, which is much more apparent for large $k$ as while simple analysis of the sub–Gaussian parameters begins to diverge from the lower bound. Figure 4 plots only the theoretical bounds to demonstrate this more clearly.